he cost structure of a manufacturer of microchips is described in the table shown below he firms fixed costs equal $10,000 per day. Calculate the average variable cost, average fxed cost, and average total cost at each output level. (Your answers should be rounded to the nearesf ce Total Cost of Output microchips per day) Average Variable Cost ($) Average Fixed Cost ($) Average Total Cost ($) $10,000 60,000 95,000 150,000 220,000 325,000 25,000 75,000 100,000 125,000 150,000
Answer #1

TC TFC TVC AVC AFC ATC N/ 0.4 0 10,000 10,000 ON/A N/A 25,000 60,000 10,000 50,000 50,000 95,000 10,000 85,000 1.7 75,000 1,50,000 10,000 1,40,000 1.866667 0.133333 1,00,000 2,20,000 10,000 2,10,000 1,25,000 3,25,000 10,000 3,15,000 2.1 2.52 2.2 0.08 1,50,000 4,65,000 10,000 4,55,000 3.033333 0.066667 3.1

TVC=Total variable cost

TFC=Total fixed cost

TC=total cost

ATC=Average total cost

=TC/Q

AVC=Average variable cost

=TVC/Q

AFC=Average fixed cost

=TFC/Q